### v1

@@ 1 Introduction
Optimization over function spaces arises in applications ranging from density estimation to signal reconstruction. We consider the general problem of optimizing an objective $\mathcal{L}:\mathcal{C}\to\mathbb{R}$ over continuous functions $f:\mathcal{D}\to\mathbb{R}$ defined on an $N$-dimensional domain $\mathcal{D}\subseteq\mathbb{R}^{N}$:
@@
where $\mathcal{C}\subseteq\mathcal{F}:=\left\{f:\mathcal{D}\to\mathbb{R}\mid\text{$f$ is continuous}\right\}$ is a constraint set.
@@
where the tensor $X\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$ represents the function $f$ evaluated on a set of grid points. The tensor has $N$ modes, matching the dimension of $\mathcal{D}$, with $I_{n}$ grid points along the $n$th dimension. The discretized objective $\tilde{\mathcal{L}}$ and constraint set $\tilde{\mathcal{C}}$ correspond to their continuous counterparts.
Any discretization raises the question, How fine should be the discretization? How large should be the grid dimensions $I_{1},\dots,I_{N}$? When we have discretization choices to make, we favor finer grids for accuracy. When data arrives pre-discretized, we prefer to incorporate all available information. In either case, fine discretization (large $I_{n}$) is desirable. Fine discretizations, however, carry computational costs: most operations are slower and consume more memory as problem size grows, often superlinearly, such as with matrix-matrix multiplication [strassen_gaussian_1969]. We propose a multiresolution approach that achieves faster convergence on fine-scale problems. Our method is inspired by wavelets [benedetto_wavelets_1993] and multigrid [trottenberg_multigrid_2001] methods, and optimizes $f$ over progressively finer discretizations.
@@ 1.1 Related Work
Many approaches exist for optimizing over function spaces beyond simple discretizations. One class of methods restricts $\mathcal{C}$ to a Hilbert space with basis functions $\{f_{k}\}$ and solves for optimal weights $\{a_{k}\}$ in the representation $f(v)=\sum_{k}a_{k}f_{k}(v)$. Examples include Fourier features, wavelets [benedetto_wavelets_1993], and Legendre polynomials [trefethen_approximation_2019]. The calculus of variations provides an analytical framework for optimizing functionals, typically integrals involving the input function [gelfand_calculus_2000].
@@
Finite parametrization offers another approach, exemplified by finding the mean $\mu$ and variance $\sigma^{2}$ in the family of normal distributions $f(v)=(2\pi\sigma)^{-1/2}\exp(-(v-\mu)^{2}/(2\sigma^{2}))$ [vershynin_HighDimensionalProbability_2018], or optimizing over weights $\{W_{k}\}$ and biases $\{b_{l}\}$ in fully connected neural networks [gurney_introduction_2018]. Non-parametric methods such as kernel density estimation [chen_tutorial_2017] and Gaussian mixture models [yan_learning_2023] provide yet another alternative. Finally, convex duality techniques transform infinite-dimensional problems into finite-dimensional counterparts [chuna_dual_2025].
@@
We focus on uniformly spaced discretizations: the approach is simple, used across a variety of settings, easy to implement and analyze, and sufficient in practice and theory for Lipschitz functions. More sophisticated discretization schemes have been well studied in differential equations [stetter_analysis_1973], probability density estimation [chakraborty_generating_2015], machine learning [ramirez_gallego_data_2016], and approximation theory [trefethen_approximation_2019]. These schemes concentrate the grid where $f$ exhibits larger variations (for differentiable $f$, where the gradient magnitude is larger) or near domain endpoints. Many results in this paper extend to such alternative discretization schemes.
@@
Our method synthesizes ideas from wavelets [benedetto_wavelets_1993], multigrid methods for partial differential equations [trottenberg_multigrid_2001], and multiresolution matrix factorization [kondor_multiresolution_2014]. At each scale $s$, we solve a distinct optimization problem ($\mathrm{P_{s}}$) that discretizes the continuous problem (P). Both the tensor dimension $X_{s}\in\mathbb{R}^{I_{s}^{1}\times\dots\times I_{s}^{N}}$ and the objective function $\tilde{\mathcal{L}}_{s}$ adapt to the scale: a coarse-scale objective involves fewer summands, different grid points, and modified constraints. This differs from approaches that merely subsample variables while keeping the objective structure fixed. By re-discretizing both the domain and the objective at each scale, our framework ensures that each subproblem $P_{s}$ properly approximates the continuous problem, not just its reduced fine-scale version.
@@
Each resolution level provides a warm start for the next finer scale, typically with a smaller error than random initialization. Warm-start strategies appear throughout numerical optimization, from stochastic gradient descent [loshchilov_sgdr_2017] to first-order methods [adcock_restarts_2025].
@@
Our coarse-scale optimization updates only a subset of coordinates (the free variables present at the current resolution) while finer-scale variables remain implicit. This connects to coordinate-wise optimization methods, including greedy descent [dhillon_nearest_2011], coordinate descent [nesterov_efficiency_2012], and block coordinate decent [xu_BlockCoordinateDescent_2013].
@@ 1.2 Contributions
We develop multiscale methods for optimizing over spaces of Lipschitz functions and establish the first theoretical framework for multiscale optimization in this setting. We propose two variants: a greedy approach that re-optimizes all variables at each scale, and a lazy approach that freezes coarse-scale values and optimizes only newly interpolated points. Both variants achieve faster convergence to tighter error bounds than gradient descent on a single fine discretization. The analysis extends to any optimization algorithm with iterate convergence at a fixed rate. We further develop constraint modification techniques that preserve feasibility across scales. Numerical tests illustrate the effectiveness of the multiscale approach on synthetic and realworld data for a distribution de-mixing problem.
@@ 1.3 Reproducible research
The data files and scripts used to generate the numerical results presented in this paper are available from the GitHub repository [Richardson_multiscale_paper_code].
@@ 2.1 Motivating Example
where the normalized Legendre polynomials are
$a_{m}(t)=\sqrt{\frac{2m+1}{2}}\sum_{k=0}^{m}\binom{m}{k}\binom{m+k}{k}\left(\frac{t-1}{2}\right)^{k}.$
The regularization term $\left\lVert p^{\prime}\right\rVert_{2}^{2}=\int_{-1}^{1}(p^{\prime}(t))^{2}\,dt$ encourages smoothness in the recovered density. We discretize this continuous problem on a uniform grid to obtain
@@
$x_{1}{[i]}=p(t_{1}{[i]})\Delta t_{1},\qquad\text{with}\qquad t_{1}[i]=-1+2\frac{i-1}{I_{1}-1}.$
@@
$A_{1}[m,i]=a_{m}(t_{1}[i])\quad\text{and}\quad G_{1}=\frac{1}{\Delta t_{1}^{3}}\begin{bmatrix}1&-1&&&\\ -1&2&-1&&\\ &\ddots&\ddots&\ddots&\\ &&-1&2&-1\\ &&&-1&1\end{bmatrix}$.
@@
Figure 1: Comparison of multiscale vs single-scale approach for the motivating example from Section˜2.1. Dots represent the median total time in milliseconds over $100$ trials; shaded regions represent the $5$th and $95$th percentile times. The multiscale approach improves algorithm speed as problem size grows. Details are provided in Section˜2.3.
@@ 2.2 Method Overview
Algorithm 1 High-level multiscale method
[…]
1:  Solve $P_{S}$ with initial point $X_{S}^{0}$ to obtain $X_{S}$.
2:for scales $s=S-1,S-2,\dots,1$ do
3:   Compute $\underline{X}_{s+1}$ by interpolating $X_{s+1}$.
4:   Solve $P_{s}$ with initial point $X_{s}^{0}=\underline{X}_{s+1}$ to obtain $X_{s}$.
5:end for
6:  Construct piecewise linear approximation $\hat{f}_{X_{1}}$ from $X_{1}$.
7:return $\hat{f}_{X_{1}}$, an approximate solution to $P$.
@@
where $X_{s}$ is a real-valued $N$-mode tensor with dimension $(I_{s}^{1}\times\dots\times I_{s}^{N})$, and $\tilde{\mathcal{L}}_{s}$ and $\tilde{\mathcal{C}}_{s}$ denote the objective and constraint set discretized at scale $s$.
@@ 2.3 Motivating Example: Numerics
To emphasise the concepts behind multiscale, we compare an instance of the multiscale method (see Algorithm˜2) against single scale optimization on the motivating example from Section˜2.1. Figure˜1 illustrates the improvement in computation time versus problem size achieved by the multiscale approach. Additional numerics and details are provided in Section˜7.
We compare total runtime for $S=3,4,\dots,12$, which yields finest discretizations with $I_{1}=2^{S}+1$ points. We initialize by uniformly discretizing the polynomial $p(t)=-2.625t^{4}-1.35t^{3}+2.4t^{2}+1.35t+0.225$ on [1,1][-1,1] to obtain the initial density approximation $x_{S}[i]=f(t_{S}[i])\Delta t_{S}$ at the coarsest scale.
At each scale $s>1$, we perform one ($K_{s}=1$) iteration; at the finest scale ($s=1$), and we iterate until the objective is within $5\%$ of its optimal value. We generate Legendre measurements for $m=1,\dots,5$ using the measurement operator $A_{1}$ from Section˜2.1 and add $5\%$ Gaussian noise to obtain $y$. The regularization parameter $\lambda=10^{-4}$ balances data fidelity against smoothness.
Figure 2: Typical progression of iterates $x_{s}^{k}$ using single scale (top row) and multiscale (bottom row) approachs for the motivating example from Section˜2.1. The clean ground truth is plotted behind in the thick dashed lined.
In Fig.˜2, we show the typical convergence of iterates $x_{s}^{k}$ for single scale and multiscale approaches. We run the motivating example from Section˜2.1 with $2^{10}+1$ points at the finest scale $s=1$. The top-left plot shows every 25th iterate $k$ starting with the initialization at $k=0$ and ending at $k=250$ when running projected gradient descent at the finest scale ($s=1$). The top-right plot shows every 250th iterate $k$, ending with the final iterate $k=4988$ which satisfies the stopping criteria $\tilde{\mathcal{L}}_{1}(x_{1}^{k})<0.00024$. The bottom-left plot shows $x_{s}^{k}$ at each scale from the coarsest scale $s=10$ to the second finest $s=2$, and the iterate at the finest scale $s=1$ after $k=25$ iterations. We can make a fair comparison between $x_{1}^{25}$ on the top-left and bottom-left plots since they are both the result of running $k=25$ iterations of projected gradient descent at the finest scale. Using multiscale results in a smoother iterate since $x_{1}^{0}$ was constructed from the chain of interpolations, rather than the completely random initialization used by the single scale approach. Because of this, the bottom-left plot highlights multiscale only needs $k=387$ iterations at the finest scale to also achieve the same stopping criteria $\tilde{\mathcal{L}}_{1}(x_{1}^{k})<0.00024$.
@@
Figure 3: Typical loss convergence using single scale (dashed blue) and multiscale (solid orange) approachs for the motivating example from Section˜2.1. The left plot shows the first millisecond, and the right plot shows the first 20 milliseconds. The left plot also highlights in grey regions that are not from projected gradient descent steps. These are the result of additional overhead with the multiscale method such as interpolation, allocations, and extra function calls.
In Fig.˜3, we show the typical loss convergence for the single scale and multiscale methods. These are shown as a function of time rather than number of iterations. Comparing iterations at different scales is not fair since we expect one projected gradient step to be cheaper at coarser scales. We also want to highlight the minimal effect of interpolation, allocations, and additional function calls. The loss $\tilde{\mathcal{L}}_{s}(x_{s}^{k})$ for the multiscale method is calculated at the respective scale $s$ of the iterate, whereas the single scale only calculates the the loss $\tilde{\mathcal{L}}_{1}(x_{1}^{k})$ at the finest scale $s=1$. This explains why the loss fluctuates up and down during interpolations for coarser scales $s$ when the approximation $\tilde{\mathcal{L}}_{s}$ for $\mathcal{L}$ is worse. The coarsest scales also take longer than expected because of the additional function call overhead. Despite these drawbacks, multiscale more than makes up for these by reducing the loss significantly faster once iterating on the finer scales. This is make clear by the right plot in Fig.˜3.
It could be argued that for optimal convergence rates with a first order method, we should not be using a fixed stepsize based on the global Lipschitz constant [nesterov-smooth-2018]. Using accelerated methods would almost certainly accelerate convergence and not require almost $5000$ iterations. But any sort of acceleration could also be used to speed up convergence for the multiscale method which already has a leg-up since the fine scale iterations start closer to the solution.
@@ 2.4 Descritization, Coarsening, and Interpolation
Consider a nested sequence of discretizations
(4) $\mathcal{D}_{S}\subset\mathcal{D}_{S-1}\subset\dots\subset\mathcal{D}_{1}\subset\mathcal{D}\subseteq\mathbb{R}^{N}$}
of the continuous domain $\mathcal{D}$. Scale $s=S$ represents the coarsest grid. Each discretization $\mathcal{D}$ contains $\left\lvert\mathcal{D}_{s}\right\rvert=\prod_{n=1}^{N}I_{s}^{n}$ points. The tensor $T$ with dimension $(I_{s}^{1}\times\dots\times I_{s}^{N})$ contains the discretization points from $\mathcal{D}_{s}$. The optimization variable $X_{s}$ represents the function values at these points:
$X_{s}[i_{1},\dots,i_{N}]=f(T_{s}[i_{1},\dots,i_{N}]).$
Unless otherwise stated, we restrict the remainder of the paper to scalar functions $f:[\ell,u]\to\mathbb{R}$ on the 1-dimensional domain $\mathcal{D}=[\ell,u]$ with uniform discretization
$x_{s}[i]=f(t_{s}[i]),\qquad\text{with}\qquad t_{s}[i]=\ell+\frac{i-1}{I_{s}-1}(u-\ell).$
@@
$\underline{x}[i]=\begin{cases}x[\frac{i+1}{2}]&\text{if $i$ is odd,}\\ \frac{1}{2}\left(x[\frac{i}{2}]+x[\frac{i}{2}+1]\right)&\text{if $i$ is even.}\end{cases}$
@@
Figure 4: Example uniform discretization of an interval with $S=3$ scales. Each scale $s$ has $2^{S-s+1}+1$ points. The free variables $x_{(s)}$ (magenta squares) resemble a “W-cycle” from multigrid [trottenberg_multigrid_2001].
@@
$\hat{f}_{x}(t)=\begin{cases}\frac{{x[2]-x[1]}}{t_{2}-t_{1}}(t-t_{1})+x[1]&\text{if $t_{1}\leq t\leq t_{2}$,}\\ \qquad\qquad\vdots\\ \frac{{x[I]-x[I-1]}}{t_{I}-t_{I-1}}(t-t_{I-1})+x[I-1]&\text{if $t_{I-1}\leq t\leq t_{I}$.}\end{cases}$
@@ 2.5 Greedy and Lazy MultiScale Algorithms
We present two variants of the multiscale method outlined in Section˜2.2 for solving Eq.˜P: the greedy variant (Algorithm˜2), which optimizes all grid points at each scale, and the lazy variant (Algorithm˜3), which optimizes only the newly interpolated points at each finer scale. Both variants apply an update rule $x^{k+1}=U(x^{k})$ at each scale, and our convergence analysis holds for any such update rule that achieves q-linear iterate convergence.
Figure 5: One hypothetical round of multiscale optimization. In all four subplots, the solid blue curve represents the solution to the continuous problem. The orange dots represent a) a random coarse initialization with four points, b) the solution at the coarse scale, c) the interpolation of the coarse solution to the fine grid, and d) the solution at the fine scale.
@@
for any feasible initialization $x^{0}\in\tilde{\mathcal{C}}$ and some minimizer $x^{*}$ of $\tilde{\mathcal{L}}(x)$ over $\tilde{\mathcal{C}}$ that may depend on $x^{0}$.
Figure˜5 illustrates one round of multiscale optimization. Suppose the solution to the continuous optimization problem is
$f(t)=-84t^{4}+146.4t^{3}-74.4t^{2}+12t$
on the interval $0\leq t\leq 1$. The figure displays the multiscale flow: initialize at the coarse scale with five points, optimize at the coarse scale, interpolate to the fine grid (adding four points), and optimize at the fine scale. In this particular example, the greedy and lazy variants share this structure but differ in the final optimization: greedy optimizes all nine fine-scale points, while lazy optimizes only the four newly interpolated points.
Our software implementation, described in Section˜7, uses projected gradient descent as the update rule, which achieves $q$-linear iterate convergence for smooth and strongly convex objectives $\tilde{\mathcal{L}}(x)$ with closed and convex constraints $\tilde{\mathcal{C}}$ [beck_first-order_2017, Sec. 10.2]. This update rule takes the form
@@
denotes the projection of $x$ onto $\tilde{\mathcal{C}}$. The standard analysis of projected gradient descent gives the following $q$-linear rate for strongly-convex objectives.
Lemma 2.7 (Descent Lemma [nesterov-smooth-2018, Theorem 2.1.15]).
Projected gradient descent with stepsize $\alpha$ has iterate convergence with rate
$q(\alpha)=\sqrt{1-\frac{2\alpha\mathcal{S}_{\tilde{\mathcal{L}}}\mu_{\tilde{\mathcal{L}}}}{\mathcal{S}_{\tilde{\mathcal{L}}}+\mu_{\tilde{\mathcal{L}}}}}$.
@@
$q(\alpha)=\frac{c-1}{c+1}$
for a condition number of $c=\mathcal{S}_{\tilde{\mathcal{L}}}/\mu_{\tilde{\mathcal{L}}}$.
@@
Algorithm 2 Greedy multiscale method
@@
[…]
3:for scales $s=S-1,S-2,\dots,1$ do
[…]
7:end for
[…]
The lazy variant freezes previously computed entries of $x$ and adjusts only the newly interpolated values $x_{(s)}$ (Definition˜2.3) during the algorithm update step. Figure˜4 illustrates these free variables. Algorithm˜3 implements this approach, differing from Algorithm˜2 only in lines 1, 2, and 6.
Algorithm 3 Lazy multiscale method
1:  Randomly initialize $x_{(s)}^{0}$ (Definition˜2.3) at the coarsest scale $s=S$.

### v2

@@ 1 Introduction
Optimization over function spaces arises in applications ranging from density estimation to signal reconstruction. We consider the general problem of optimizing an objective $\mathcal{L}:\mathcal{C}\to\mathbb{R}$ over continuous functions $f:\mathcal{D}\to\mathbb{R}$ defined on an $1$-dimensional domain $\mathcal{D}\subseteq\mathbb{R}$:
@@
where $\mathcal{C}\subseteq\mathcal{F}:=\left\{f:\mathcal{D}\to\mathbb{R}\mid f\ \text{is continuous}\right\}$ is a constraint set.
@@
where the vector $x\in\mathbb{R}^{I}$ represents the function $f$ evaluated on a set of grid points. The discretized objective $\tilde{\mathcal{L}}$ and constraint set $\tilde{\mathcal{C}}$ correspond to their continuous counterparts.
Any discretization raises the question; how fine should be the discretization? When we have discretization choices to make, we often favor finer grids for accuracy. When data arrives pre-discretized, we prefer to incorporate all available information. Fine discretizations, however, carry computational costs: most operations are slower and consume more memory as problem size grows, often superlinearly, such as with matrix-matrix multiplication [23]. We propose a multiresolution approach that achieves faster convergence on fine-scale problems. Our method is inspired by wavelets [3] and multigrid [25] methods, and optimizes $f$ over progressively finer discretizations.
@@ 1.1 Related Work
Many approaches exist for optimizing over function spaces beyond simple discretizations. One class of methods restricts $\mathcal{C}$ to a Hilbert space with basis functions $\{f_{k}\}$ and solves for optimal weights $\{a_{k}\}$ in the representation $f(v)=\sum_{k}a_{k}f_{k}(v)$. Examples include Fourier features, wavelets [3], and Legendre polynomials [24]. The calculus of variations provides an analytical framework for optimizing functionals, typically integrals involving the input function [9].
@@
Finite parametrization offers another approach, exemplified by finding the mean $\mu$ and variance $\sigma^{2}$ in the family of normal distributions [26], or optimizing over weights $\{W_{k}\}$ and biases $\{b_{l}\}$ in fully connected neural networks [12]. Non-parametric methods such as kernel density estimation [5] and Gaussian mixture models [28] provide yet another alternative. Finally, convex duality techniques transform infinite-dimensional problems into finite-dimensional counterparts [6].
@@
We focus on uniformly spaced discretizations: the approach is simple, used across a variety of settings, easy to implement and analyze, and sufficient in practice and theory for Lipschitz functions. More sophisticated discretization schemes have been well studied in differential equations [22], probability density estimation [4], machine learning [19], and approximation theory [24]. These schemes concentrate the grid where $f$ exhibits larger variations (for differentiable $f$, where the gradient magnitude is larger) or near domain endpoints. Many results in this paper extend to such alternative discretization schemes.
@@
Our method synthesizes ideas from wavelets [3], multigrid methods for partial differential equations [25], and multiresolution matrix factorization [10, 14]. At each scale $s$, we solve a distinct optimization problem ($\mathrm{P_{s}}$) that discretizes the continuous problem (P). Both the vector’s dimension $I$ and the objective function $\tilde{\mathcal{L}}_{s}$ adapt to the scale: a coarse-scale objective involves fewer summands, different grid points, and modified constraints. This differs from approaches that merely subsample variables while keeping the objective structure fixed. By re-discretizing both the domain and the objective at each scale, our framework ensures that each subproblem $P_{s}$ properly approximates the continuous problem, not just its reduced fine-scale version. This contrasts more direct translations of multigrid for smooth optimization problems like MGProx [2] which uses a V-cycle to start with the finest scale problem and transfer information to the coarsest scale before working back to the finest scale problem. Our approach starts with the coarsest scale and only works towards the finest scale problem. We also study a more general problem that does not rely on a specific inner algorithm.
@@
Each resolution level provides a warm start for the next finer scale, typically with a smaller error than random initialization. Warm-start strategies appear throughout numerical optimization, from stochastic gradient descent [16] to first-order methods [1].
@@
Our coarse-scale optimization updates only a subset of coordinates (the free variables present at the current resolution) while finer-scale variables remain implicit. This connects to coordinate-wise optimization methods, including greedy descent [8], coordinate descent [17], and block coordinate decent [27].
@@ 1.2 Contributions
We develop multiscale methods for optimizing over spaces of Lipschitz functions and establish the first theoretical framework for multiscale optimization in this setting. We propose two variants: a greedy approach that re-optimizes all variables at each scale (Algorithms˜1 and 4.2), and a lazy approach that freezes coarse-scale values and optimizes only newly interpolated points (Algorithms˜1 and 4.3). Both variants achieve faster convergence to tighter error bounds than gradient descent on a single fine discretization (Corollaries˜6.8 and 6.14). The analysis extends to any optimization algorithm with iterate convergence at a fixed rate (Definition˜2.6). We further develop constraint modification techniques that preserve feasibility across scales (Section˜3), and bound the error between our algorithm output and the solution to the continuous problem (Section˜5). Numerical tests illustrate the effectiveness of the multiscale approach on synthetic (Sections˜7.2 and 2.3) and realworld (Section˜7.3) data for a distribution de-mixing problem. We make clear that although the theoretical analysis is given to solve $1$-dimensional problems, the general principle extends to higher dimensions where the objective variable becomes one or more tensors (Section˜7).
@@ 1.3 Reproducible research
The data files and scripts used to generate the numerical results presented in this paper are available from the GitHub repository [21].
@@ 2.1 Motivating Example
with normalized Legendre polynomials $a_{m}(t)=(\tfrac{2m+1}{2})^{1/2}\sum_{k=0}^{m}\binom{m}{k}\binom{m+k}{k}\left(\frac{t-1}{2}\right)^{k}.$ The regularization term $\left\lVert p^{\prime}\right\rVert_{2}^{2}=\int_{-1}^{1}(p^{\prime}(t))^{2}\,dt$ encourages smoothness in the recovered density. We discretize this continuous problem on a uniform grid to obtain
@@
$x_{1}{[i]}=p(t_{1}{[i]})\Delta t_{1},\qquad\text{with}\qquad t_{1}[i]=-1+2\tfrac{i-1}{I_{1}-1}.$
@@
[…]
@@
Figure 1: Comparison of multiscale vs single-scale approach for the motivating example from Section˜2.1. Dots represent the median total time in milliseconds over $100$ trials; shaded regions represent the $5$th and $95$th percentile times. The multiscale approach improves algorithm speed over a single scaled approach once the problem size is large enough. Details are provided in Section˜2.3.
@@ 2.2 Method Overview
where $x_{s}\in\mathbb{R}^{I_{s}}$, and $\tilde{\mathcal{L}}_{s}$ and $\tilde{\mathcal{C}}_{s}$ denote the discretized objective and constraint set.
@@ 2.3 Motivating Example: Numerics
To emphasise the concepts behind multiscale, we compare an instance of the multiscale method (see Algorithm˜1) against single scale optimization on the motivating example from Section˜2.1. Additional numerics and details are provided in Sections˜7 and LABEL:sec-motivating-example-details.
Figure˜1 illustrates the improvement in computation time versus problem size achieved by the multiscale approach.
Figure 2: Typical progression of iterates $x_{s}^{k}$ for scale $s$ and iteration $k$ using single scale (top row) and multiscale (bottom row) approaches for the motivating example from Section˜2.1. The clean ground truth is plotted behind in the thick dashed lined.
Figure˜2 shows the typical convergence of iterates $x_{s}^{k}$ for single scale and multiscale approaches. And Fig.˜3 shows the typical loss convergence for the single scale and multiscale methods.
@@
Figure 3: Typical loss convergence using single scale (dashed blue) and multiscale (solid orange) approaches for the motivating example from Section˜2.1. The left plot shows the first millisecond, and the right plot shows the first 20 milliseconds. The left plot also highlights in grey regions that are not from projected gradient descent steps. These are the result of minimal overhead with the multiscale method such as interpolation, allocations, and extra function calls.
@@ 2.4 Descritization, Coarsening, and Interpolation
Unless otherwise stated, we restrict the focus of this paper to scalar functions $f:[\ell,u]\to\mathbb{R}$ on the 1-dimensional domain $\mathcal{D}=[\ell,u]$ with uniform discretization
$x_{s}[i]=f(t_{s}[i]),\qquad\text{with}\qquad t_{s}[i]=\ell+\tfrac{i-1}{I_{s}-1}(u-\ell).$
@@
$\underline{x}[i]=\begin{cases}x[\tfrac{i+1}{2}]&\text{if $i$ is odd,}\\ \tfrac{1}{2}\left(x[\tfrac{i}{2}]+x[\tfrac{i}{2}+1]\right)&\text{if $i$ is even.}\end{cases}$
@@
Figure 4: Discretization of an interval with $S=3$ scales. Each scale $s$ has $2^{S-s+1}+1$ points. Starting at scale $s=3$, newly added points at each scale (the free variables $x_{(s)}[i]$) are shown in purple squares, while existing points are shown in teal circles.
@@
$\hat{f}_{x}(t)=\tfrac{{x[i+1]-x[i]}}{t_{i+1}-t_{i}}(t-t_{i})+x[i]\quad\text{where}\quad t_{i}\leq t\leq t_{i+1}\quad\text{for}\quad i=1,\dots,I.$
@@ 2.5 Greedy and Lazy Multiscale Algorithms
We present two variants of the multiscale method (Algorithm˜1) outlined in Section˜2.2 for solving Eq.˜P: the greedy variant, which optimizes all grid points at each scale, and the lazy variant, which optimizes only the newly interpolated points at each finer scale. Both variants apply an update rule $x^{k+1}=U(x^{k})$ at each scale, and our convergence analysis holds for any such update rule that achieves $q$-linear iterate convergence.
@@
for any initialization $x^{0}\in\tilde{\mathcal{C}}$ and minimizer $x^{*}$ of $\tilde{\mathcal{L}}(x)$ over $\tilde{\mathcal{C}}$ that may depend on $x^{0}$.
Our software implementation (Section˜7), uses projected gradient descent as the update rule, which is described by
@@
Lemma 2.7 (Descent Lemma [18, Theorem 2.1.15]).
@@
$q(\alpha)=\sqrt{1-\tfrac{2\alpha\mathcal{S}_{\tilde{\mathcal{L}}}\mu_{\tilde{\mathcal{L}}}}{\mathcal{S}_{\tilde{\mathcal{L}}}+\mu_{\tilde{\mathcal{L}}}}}$.
[…]
@@
Algorithm 1 The Multiscale Method: step 6 specifies Greedy or Lazy versions